A juice drink manufacturer has found mixing pineapple juice concentrate and fresh squeezed orange juice is the most cost effective way to produce its breakfast beverage. The pineapple juice comes in two concentrated forms, 24% and 60% pineapple juice, which must be combined into a solution that can be mixed with the orange juice. If 50 gallons of 24% juice is used, how much of the 60% just must be used to obtain a 50% pineapple juice solution?
Answer by mananth(16946)

percent ---------------- quantity
24 ---------------- 50
60 ---------------- x
50 ---------------- 50+x
...
24*50+60 *x=50(50 +x)
1200+60x =2500+50 x
60x-50x= 2500-1200
10x=1300
/10
x=130 gallons of 60%